I have two question
1) A 25.0mL sample of HClO4 solution requires 34.6mL of 0.100 M NaOH for complete neutralization. what was the concentration of the HClO4 solution?
HClO4(aq)+NaOH(aq)-H2O(1)+NaClO4(aq)

2) A 683 mL NaOH solution is diluted to a volume of 1.75 L and has a final concentration of 5.00 M what was the initial concentration?
1
Expert's answer
2019-05-06T05:05:43-0400

1.A 25.0mL sample of HClO4 solution requires 34.6mL of 0.100 M NaOH for complete neutralization. what was the concentration of the HClO4 solution? 

Solution:

HClO4(aq)+NaOH(aq)=H2O(1)+NaClO4(aq) 

n(NaOH)=0.1*0.0346=0.00346 mol

n(NaOH)=n(HClO4).It is with reaction.

CM(HClO4)=0.00346/0.025=0.14 M

Answer: CM=0.14 M.

2.A 683 mL NaOH solution is diluted to a volume of 1.75 L and has a final concentration of 5.00 M what was the initial concentration?

Solution:

n(NaOH)=5*1.75=8.75 mol;

CM(NaOH)=8.75*0.683=5.98 M

Answer: CM(NaOH)=5.98 M.